AB INITIO CHARACTERISATION OF THE STABLE CONFORMERS OF C_{4} SUGARS: ERYTHROSE, ERYTHRULOSE AND THREOSE

\underline{J.R. AVILES-MORENO},D. PETITPREZ, and T.R. HUET, Laboratoire PhLAM, UMR 8523 CNRS-Universite de Lille 1, F-59655 Villeneuve d’Ascq Cedex, France.

The importance of sugars (C_{n}H_{2}nO_{n}) in biology and exobiology motivates this study (i. e. the detection of glycolaldehyde (C_{2}H_{4}O_{2}) in Sgr2(N-LMH) cloud by J.M. HOLLIS et al. Astrophys. J. Lett. 540(2000)L107 and the microwave study of glyceraldehyde (C_{3}H_{6}O_{3}) by F.J.LOVAS et al. J.Mol.Spect. 222(2003),263). To the best of our knowlege, there is no experimental characterisation of C_{4} sugars in the gas phase. Therefore we have performed ab initio calculations for the C_{4} sugars in order to help further identifications. We have calculated equilibrium rotational constants, electric dipole moments and relative energies of different conformations. The expected accuracy is 0.5% on rotational constants and 1.0% kJ/mol on relative energies. Concerning erythrose, 14 conformations have been found at the B3LYP/6-311++G(2df,p) level of the theory. The relative energy of all conformations has been optimized at the G3MP2B3 level of the theory. A few conformations are found within a small range of energy, under 5 kJ/mol. These most stable structures are stabilized by intra-molecular hydrogen bonds. Results on rotational constants and electric dipole moments performed at the B3LYP and the MP2 levels of the theory with 6-311++G(2df,p) basis sets suggest that a detection in the microwave region could be possible, using for example FTMW spectroscopy coupled to a molecular beam. Calculations for erythrulose and threose are in progress. This work is supported by grants IDRIS number 041715 and 051715.